# LLMs are Also Effective Embedding Models: An In-depth Overview

### v1

@@ 6 Advanced Embeddings
Beyond regular text embeddings, advanced techniques have been developed to handle longer texts, multiple languages, cross-modal data, and programming code. This section delves into these advanced embedding techniques, highlighting their methodologies and applications.
6.1 Multi-lingual Embedding
@@
Multi-lingual embeddings are designed to represent text from different languages in a shared vector space. These embeddings are useful for various tasks such as cross-lingual information retrieval, machine translation, and multi-lingual sentiment analysis. The development of monolingual language models has substantially progressed in learning embeddings enriched with contextual information across a range of specific languages. Early pretrained LMs like BERT  (Devlin et al., 2018) and GPT  (Radford et al., 2018) learn fine-grained contextual monolingual representations through masked language modeling (MLM) and next-token prediction (NTP). Built on the foundation of monolingual LMs, a series of advanced works have been proposed to learn universal multi-lingual embeddings. We categorize these approaches based on their alignment techniques into two types: implicit alignment and explicit alignment.
6.1.1 Implicit alignment.
@@
Jina-embeddings-v3 (Sturua et al., 2024) is a multilingual embedding model that employs task-specific Low-Rank Adaptation (LoRA) adapters to enhance performance across various downstream tasks. The model supports long-context retrieval with sequences up to 8192 tokens by integrating RoPE (Su et al., 2024). Additionally, jina-embeddings-v3 incorporates Matryoshka Representation Learning (Kusupati et al., 2024) to allow flexible reduction of embedding dimensions without compromising performance, enabling embeddings to be truncated to as low as 32 dimensions. By addressing common retrieval failures using synthetic data and leveraging instruction tuning (Wei et al., 2021; Su et al., 2023a), the model achieves state-of-the-art results on the MTEB benchmark (Muennighoff et al., 2022), outperforming proprietary embeddings from OpenAI and Cohere on English tasks and surpassing multilingual-e5-large-instruct (Wang et al., 2022b) across all multilingual tasks.
6.1.2 Explicit Alignment.
@@
Schmidt et al. (2024) aligned the multilingual representations, which were extracted by machine translation encoder, into the semantic embedding space of LLMs via self-distillation. This integration enables the LLMs to perform zero-shot ability to any language supported by the machine translation encoder. Vasilyev et al. (2024) employed a methodology that exclusively fine-tunes the query encoder while keeping the text encoder frozen on an english-only dataset, finding that this approach not only preserves but significantly enhances the multilingual embedding capabilities of the model. Based on the assumption that monolingual embeddings are well-structured, Park et al. (2024) proposed to distill alignment information from the monolingual similarity matrix into cross-modal embeddings to guide the cross-lingual alignment process. This method resolved issues in prior contrastive learning approaches that treated non-exact translation pairs as negative samples, which disrupted the monolingual embedding space. In addition, to further assess the capabilities of multilingual language models, Yoon et al. (2024) presented MINERS, a multilingual, multi-task, tuning-free benchmark specifically designed to evaluate semantic retrieval tasks including bitext mining and classification via retrieval-augmented contexts.
6.2 Long Context Embedding
@@
The ability to effectively capture and utilize long context information has become increasingly vital. Traditional embedding techniques, which typically focus on limited length of text, often struggle to maintain coherence and capture the full semantic meaning when applied to texts with extended contexts. Long context embeddings are designed to address these challenges, enabling models to handle extended sequences of text while preserving contextual integrity and meaning. A popular method for extending the length of context is training the embedding models with the long-context backbone, which can be obtained either by using the existing model or by pre-training with long inputs from scratch (Wang et al., 2024a; Nussbaum et al., 2024; Günther et al., 2023; Chen et al., 2024). Typically, Zhang et al. (2024b) introduced a text encoder enhanced with RoPE (Su et al., 2024) and unpadding portes2024mosaicbert, which is pre-trained by masked language model based on much longer (e.g. 8K tokens) and multilingual context with a two-stage curriculum. Based on the text encoder, they construct the embedding model through contrastive pre-training and finetuning utilizing InfoNCE as the loss function.
@@
However, fine-tuning LLMs may result in high computational costs, where there are more and more efforts are proposed to extend the contexts of LLM embedding models through the plug-and-play methods (Ratner et al., 2022; Jin et al., 2024; Wang et al., 2024b). Luo et al. (2024a) proposed a chunking-free architecture to process long context, adding a special token <LMK> at the end of each sentence in the long context to capture the coherent semantic. During training, this work introduced the positional weight and utilizes multi-stage training to achieve a superior performance. Shao et al. (2024) proposed a plug-in module called the extensible embedder to process long contexts, which were partitioned into multiple chunks. Each chunk was embedded and down-scaled by the extensible embedder as a compact representation. This work trained the extensible embedder through two-stream auto-regression while keeping the downstream LLM’s parameters fixed, thereby preserving the LLM’s original capabilities. Furthermore, Zhu et al. (2024) explored extensive plug-and-play strategies to extend existing embedding models to long contexts, including parallel context windows, grouped positions, recurrent positions, linear position interpolation, and more. This work also proposed the newly constructed LONGEMBED benchmark for long context retrieval evaluation, which included two synthetic tasks with flexible document lengths and four real tasks featuring dispersed target information.
6.3 Cross-modal Embedding
@@
The potential of Transformers for efficiently learning cross-modal embeddings has been demonstrated by works like VATT  (Akbari et al., 2021), ViT (Dosovitskiy et al., 2020). Following the introduction of CLIP  (Radford et al., 2021), which uses language as supervision for visual models to bridge the gap between vision and language, a surge of vision-language cross-modal models such as ALBEF  (Li et al., 2021), VLMO  (Bao et al., 2021b), CoCa  (Yu et al., 2022), BLIP  (Li et al., 2022), and BEiT  (Bao et al., 2021a) emerged. However, these models have significant limitations in representing purely textual or purely visual data. VISTA  (Zhou et al., 2024) proposes a new embedding approach for universal multimodal retrieval, allowing the pretrained LMs to recognize image tokens by utilizing ViT as an image tokenizer. It undergoes two-stage training on datasets with weak labels and high-quality synthetically composed image-text datasets, achieving superior performance across a variety of multimodal retrieval tasks.
@@
With the development of LLMs, many efforts have extended LLMs to various modalities, such as image (BLIP-2  (Li et al., 2023a), MiniGPT4  (Zhu et al., 2023), Llava  (Liu et al., 2024a)), audio (AudioPaLM  (Rubenstein et al., 2023), Qwen-Audio  (Chu et al., 2023)), and video (Video-Llama  (Zhang et al., 2023a), Video-ChatGPT  (Maaz et al., 2023), Video-Llava  (Lin et al., 2023)). However, these efforts primarily focus on generating text related to these modalities. PaLM 2 DE  (Gomez et al., 2024) incorporates LLMs into dual-encoder architecture, utilizing PaLM to initialize and augment multilingual text comprehension for cross-modal retrieval tasks. This integration results in enhanced performance across an impressive range of 102 languages, even though it was trained on a limited dataset of 21 languages. Furthermore, by leveraging machine translation data, PaLM 2 DE significantly boosts its cross-lingual capabilities.
7 Discussions
@@
Table 4: Results on MTEB. The models are categorized into two groups based on their embedding dimension (Dim.) and the number of parameters (# Params.).
@@
E5-mistral-7b-instruct (Wang et al., 2024a) 4,096 7B 78.47 50.26 88.34 60.21 56.89 84.63 31.40 66.63
@@
Dense vs. Sparse Embedding. While constructing dense embeddings based on LLMs has gained significant attention, some researchers are also investigating the acquisition of sparse embeddings by leveraging lexicon-importance distributions derived from LLMs (Chen et al., 2024; Zhuang et al., 2024; Doshi et al., 2024; Nie et al., 2024). Zhuang et al. (2024) evaluated various embedding types by directly prompting or fine-tuning LLMs. Their findings reveal that in zero-shot retrieval settings, dense and sparse embeddings perform differently across different benchmarks: for instance, dense embeddings outperform sparse ones on BEIR and TREC-DL 2019, while sparse embeddings show better results on TREC-DL 2020 and MSMARCO. However, in supervised fine-tuning settings, dense retrieval significantly outperforms sparse retrieval on datasets like MSMARCO and TREC-DL 2019/2020 when using models such as Llama3-8B and Llama3-70B. Notably, combining both representation types further enhances performance. Chen et al. (2024) fine-tuned multilingual, multi-granularity text embeddings, integrating both dense and sparse retrieval models based on XLM-RoBERTa. Their results show that dense retrieval notably outperforms sparse retrieval on cross-lingual tasks such as MKQA and MIRACL, and that combining both methods also yields superior results, consistent with the findings reported in (Zhuang et al., 2024).
Figure 3: The performance trend of models with varying sizes on MTEB benchmark, based on results from Muennighoff et al. (2022).
@@
Last token vs. Mean pooling. Two common methods for obtaining embeddings from a sequence of tokens are: i) mean pooling and ii) the embedding of the last [EOS] token. The former calculates the average of token embeddings, which can potentially obscure key information from phrases. In contrast, [EOS] token embedding is susceptible to recency bias, as it heavily relies on the embedding of the final token in the sequence. Lee et al. (2024a) conducted experiments comparing several methods for obtaining embeddings from the last layer of LLMs, including the [EOS] token embedding, mean pooling, and their proposed latent attention method. The results demonstrate that mean pooling can always better than [EOS] token embedding based on average MTEB scores, regardless of whether causal attention or bidirectional attention settings are used in the LLMs. Furthermore, their latent attention method enhances embedding capabilities, yielding better results than both the [EOS] token embedding and mean pooling. This research highlights the potential for exploring more advanced designs to improve last-layer embeddings for enhanced performance.
@@
In-context Learning (ICL) for Embedding Model. ICL has emerged as a powerful paradigm for enhancing the performance of LLMs in various generative tasks. It enables LLMs to adapt dynamically to the input context by leveraging real-time examples and instructions during inference. This capability to learn from the surrounding context potentially allows LLMs to produce more accurate and context-sensitive embeddings. Recently, researchers have also explored the effectiveness of ICL for LLM-based representation models. For instance, Jiang et al. (2023b) first investigated incorporating ICL examples into prompt-based methods and demonstrated their effectiveness for STS tasks. Their approach involved constructing ICL examples by generating word-sentence pairs using ChatGPT to align words with sentence semantics from the STS-B training set, as well as utilizing word-definition pairs from the Oxford dictionary, inspired by DefSent (Tsukagoshi et al., 2021). More recently, Li et al. (2024b) incorporated few-shot examples on the query side to enhance query embeddings during supervised contrastive training. Despite these few-shot examples being randomly sampled, they still improved the model’s performance. Besides, Springer et al. (2024) proposed repeating the input twice within the context to obtain sequence embeddings, a technique that can be viewed as a specialized form of ICL, and validated its effectiveness in both training-free and supervised contrastive training settings. These studies collectively indicate that ICL is a promising strategy for enhancing representation learning with LLMs. Nevertheless, exploring more diverse and high-quality contextual information holds potential for further improving the effectiveness of this approach.
@@
Scaling Law of Embedding Models. The question of whether scaling laws hold for embedding models based on LLMS is both significant and intriguing. The effectiveness of embedding models is highly sensitive to several training factors, including model size, data size, and data quality (e.g., diversity and quality of negatives), making it difficult to isolate the effect of each factor independently. Fang et al. (2024) evaluated the quality of dense retrieval models using a contrastive entropy metric and observe a power law relationship between model performance, model size, and data size across various annotation methods and datasets. However, their experiments were conducted on relatively small-scale models, such as BERT models ranging from 0.5M to 110M parameters. Additionally, Muennighoff et al. (2022) reported the performance of models with varying sizes, such as SGPT (125M, 1.3B, 2.7B, 5.8B) and GTR (Base, Large, XL, XXL), on the MTEB benchmark shown in Figure 3. Their results show consistent performance improvements with increasing model size, even up to 5.8B parameters. Furthermore, Zhuang et al. (2024) tested the performance of various embedding types by directly prompting LLMs and found that Llama3-70B-Instruct outperformed Llama3-8B-Instruct across different embedding types. All these evidences suggest that the scaling law may also hold for embedding models based on larger-scale LMs, although more rigorous experiments are needed to confirm this.
8 Open Problems

### v2

@@ 6. Specialized Embeddings
Beyond regular text embeddings, advanced techniques have been developed to handle longer texts, multiple languages, cross-modal data, and programming code. This section delves into these specialized embedding techniques, highlighting their methodologies and applications.
6.1. Multi-lingual Embedding
@@
Multi-lingual embeddings are designed to represent text from different languages in a shared vector space. These embeddings are useful for various tasks such as cross-lingual information retrieval, machine translation, and multi-lingual sentiment analysis. The development of monolingual language models has substantially progressed in learning embeddings enriched with contextual information across a range of specific languages.
Early pretrained LMs like BERT  (Devlin et al., 2018) and GPT  (Radford et al., 2018) learn fine-grained contextual monolingual representations through masked language modeling (MLM) and next-token prediction (NTP). Built on the foundation of monolingual LMs, a series of advanced works have been proposed to learn universal multi-lingual embeddings. We categorize these approaches based on their alignment techniques into two types: implicit alignment and explicit alignment.
6.1.1. Implicit alignment.
@@
(Sturua et al., 2024) introduced jina-embeddings-v3, a multilingual embedding model employs task-specific Low-Rank Adaptation (LoRA) adapters to enhance performance across various downstream tasks. The model supports long-context retrieval with sequences up to 8192 tokens by integrating RoPE (Su et al., 2024). Additionally, jina-embeddings-v3 incorporates Matryoshka Representation Learning (Kusupati et al., 2024) to allow flexible reduction of embedding dimensions without compromising performance, enabling embeddings to be truncated to as low as 32 dimensions. By addressing common retrieval failures using synthetic data and leveraging instruction tuning (Wei et al., 2021; Su et al., 2023a), the model achieves state-of-the-art results on the MTEB benchmark (Muennighoff et al., 2022), outperforming proprietary embeddings from OpenAI and Cohere on English tasks and surpassing multilingual-e5-large-instruct (Wang et al., 2022b) across all multilingual tasks.
6.1.2. Explicit Alignment.
@@
MT-LLM  (Schmidt et al., 2024) aligns the multilingual representations, which are extracted by machine translation encoder, into the semantic embedding space of LLMs via self-distillation. This integration enables the LLMs to perform zero-shot ability to any language supported by the machine translation encoder. (Vasilyev et al., 2024) employs a methodology that exclusively fine-tunes the query encoder while keeping the text encoder frozen on an english-only dataset, finding that this approach not only preserves but significantly enhances the multilingual embedding capabilities of the model. Based on the assumption that monolingual embeddings are well-structured, (Park et al., 2024) proposes to distill alignment information from the monolingual similarity matrix into cross-modal embeddings to guide the cross-lingual alignment process. This method resolves issues in prior contrastive learning approaches that treated non-exact translation pairs as negative samples, which disrupted the monolingual embedding space. In addition, to further assess the capabilities of multilingual language models,  (Yoon et al., 2024) propose MINERS, a multilingual, multi-task, tuning-free benchmark specifically designed to evaluate semantic retrieval tasks including bitext mining and classification via retrieval-augmented contexts.
Table 3. Recent Benchmarks for Evaluating Multilingual and Language-Specific.
Benchmark Language Coverage Year
MMTEB (Enevoldsen et al., [n. d.]) 250+ languages 2025
mFollowIR (Weller et al., 2025) Persian, Chinese, Russian 2025
FaMTEB (Zinvandi et al., 2025) Persian 2025
C-MTEB (Xiao et al., 2024b) Chinese 2024
PL-MTEB (Poświata et al., 2024) Polish 2024
Scandinavian Embedding Benchmark (SEB) (Enevoldsen et al., 2024) Nordic languages 2024
ArabicMTEB (Bhatia et al., 2024) Arabic dialects, cross-lingual 2024
RusBEIR (Kovalev et al., 2025) Russian 2024
JMTEB (Li et al., 2024c) Japanese 2024
Mteb-french (Ciancone et al., 2024) French 2024
Amharic Passage Retrieval Benchmark (Amde Mekonnen et al., 2025) Amharic 2025
6.1.3. Multilingual Text Embedding Benchmark
Recent efforts have also focused on benchmarking multilingual embeddings to evaluate and advance their effectiveness. Among these, the Massive Multilingual Text Embedding Benchmark (MMTEB) (Enevoldsen et al., [n. d.]), an extension of the well-known MTEB (Muennighoff et al., 2022), provides a comprehensive evaluation across over 250 languages and 500 tasks spanning 10 categories. In addition to MMTEB, specialized multilingual benchmarks have been proposed for specific languages or language groups, such as mFollowIR (Weller et al., 2025) for instruction-following retrieval, FaMTEB (Zinvandi et al., 2025) for Persian, C-MTEB (Xiao et al., 2024b) for Chinese, PL-MTEB (Poświata et al., 2024) for Polish, the Scandinavian Embedding Benchmarks (Enevoldsen et al., 2024) covering Nordic languages, Swan and ArabicMTEB (Bhatia et al., 2024) focusing on Arabic dialects and cross-cultural embeddings, as well as JMTEB (Li et al., 2024c) for Japanese and various benchmarks targeting German (Wehrli et al., 2024), targeting French (Ciancone et al., 2024), Russian (Kovalev et al., 2025), and Amharic (Amde Mekonnen et al., 2025). These benchmarks collectively address the linguistic diversity, dialectal variations, and cultural contexts inherent in multilingual NLP, providing crucial standardized frameworks for evaluating embedding quality and facilitating progress in universal and language-specific embedding models.
6.2. Code Embedding
LLM-based code embeddings transform programming code into continuous vector representations that capture semantic, syntactic, functional properties and logic rules. These embeddings enable retrieval systems to encode code similarly to natural language, powering various types of code retrieval tasks. Traditional text embedding approaches struggle with code retrieval tasks which is largely due to a mismatch between the semantic structures of code and natural language. Therefore, it’s important to align features within and between code and natural languages. Early code embedding works like code2vec (Alon et al., 2019) and Codebert (Feng et al., 2020), adapted NLP techniques to code by treating code as text. However, these methods struggled with structural nuances (e.g., control flow, data dependencies) and multi-language support. GraphCodeBERT (Guo et al., 2020) improved this by incorporating abstract syntax trees (ASTs). Further, UniXcoder (Guo et al., 2022) unifies code representations (tokens, ASTs, comments) into a single multimodal transformer, enabling diverse code understanding and generation tasks like code search and summarization.
With the development of LLMs, some LLM-based code embedding models were proposed to alleviate the limitation.  (Li et al., 2023a) released StarCoderBase and StarCoder models, trained on large amounts of code data across more than 80 programming languages which were sourced from GitHub issues, Git commits, and Jupyter notebooks. They also used multi-query attention for efficient long-context embedding (8K tokens).  (Liu et al., 2024c) proposed a generalizable training framework which converted diverse code-related tasks into retrieval tasks, and constructed training samples from code-to-text, text-to-code, code-to-code, text-to-text and hybrid text and code tasks across multiple programming languages.  (Li et al., 2025a) introduced CodeR trained on a large-scale synthetic training dataset, which was created via a novel pipeline using LLMs for task design and sample generation. Meanwhile they employed a three-stage curriculum learning strategy to optimize training process.
6.3. Long Context Embedding
@@
The ability to effectively capture and utilize long context information has become increasingly vital. Traditional embedding techniques, which typically focus on limited length of text, often struggle to maintain coherence and capture the full semantic meaning when applied to texts with extended contexts. Long context embeddings are designed to address these challenges, enabling models to handle extended sequences of text while preserving contextual integrity and meaning. A popular method for extending the length of context is training the embedding models with the long-context backbone, which can be obtained either by using the existing model or by pre-training with long inputs from scratch (Wang et al., 2024c; Nussbaum et al., 2024; Günther et al., 2023; Chen et al., 2024). Typically, (Zhang et al., 2024b) introduced a text encoder enhanced with RoPE (Su et al., 2024) and unpadding (Portes et al., 2024), which is pre-trained by masked language model based on much longer (e.g. 8K tokens) and multilingual context with a two-stage curriculum. Based on the text encoder, they construct the embedding model through contrastive pre-training and finetuning utilizing InfoNCE as the loss function.
@@
However, fine-tuning LLMs may result in high computational costs, where there are more and more efforts are proposed to extend the contexts of LLM embedding models through the plug-and-play methods (Ratner et al., 2022; Jin et al., 2024; Wang et al., 2024a). (Luo et al., 2024a) proposes a chunking-free architecture to process long context, which adds a special token ¡LMK¿ at the end of each sentence in the long context to capture the coherent semantic. During training, this work introduces the positional weight and utilizes multi-stage training to achieve a superior performance. (Shao et al., 2024) proposes a plug-in module extensible embedder to process the long context which is partitioned into multiple chunks. Each chunk is embedded and down-scaled by extensible embedder as the compact representation. This work train extensible embedder through two-stream auto-regression with fixed downstream LLM’s parameters, which does not affect the LLM’s original capabilities. Furthermore, (Zhu et al., 2024) explores the extensive plug-and-play strategies to extend existing embedding models to long context, which includes parallel context windows, grouped positions, recurrent positions, linear position interpolation and so on. This work also proposes a newly constructed LONGEMBED benchmark for long context retrieval evaluation, which includes two synthetic tasks with flexible document length and four real tasks with featuring dispersed target information.
6.4. Cross-modal Embedding
@@
The potential of Transformers for efficiently learning cross-modal embeddings has been demonstrated by works like VATT  (Akbari et al., 2021), ViT  (Dosovitskiy et al., 2020). Following the introduction of CLIP  (Radford et al., 2021), which uses language as supervision for visual models to bridge the gap between vision and language, a surge of vision-language cross-modal models such as ALBEF  (Li et al., 2021), VLMO  (Bao et al., 2021b), CoCa  (Yu et al., 2022), BLIP  (Li et al., 2022), and BEiT  (Bao et al., 2021a) emerged. However, these models have significant limitations in representing purely textual or purely visual data. The VISTA  (Zhou et al., 2024) proposes a new embedding approach for universal multimodal retrieval, allowing the pretrained LMs to recognize image tokens by utilizing ViT as an image tokenizer. It undergoes two-stage training on datasets with weak labels and high-quality synthetically composed image-text datasets, achieving superior performance across a variety of multimodal retrieval tasks.
@@
With the development of LLMs, many efforts have extended LLMs to various modalities, such as image (BLIP-2  (Li et al., 2023c), MiniGPT4  (Zhu et al., 2023), Llava  (Liu et al., 2024b)), audio (AudioPaLM  (Rubenstein et al., 2023), Qwen-Audio  (Chu et al., 2023)), and video (Video-Llama  (Zhang et al., 2023a), Video-ChatGPT  (Maaz et al., 2023), Video-Llava  (Lin et al., 2023)). However, these efforts primarily focus on generating text related to these modalities. PaLM 2 DE  (Gomez et al., 2024) incorporates LLMs into dual-encoder architecture, utilizing PaLM to initialize and augment multilingual text comprehension for cross-modal retrieval tasks. This integration results in enhanced performance across an impressive range of 102 languages, even though it was trained on a limited dataset of 21 languages. Furthermore, by leveraging machine translation data, PaLM 2 DE significantly boosts its cross-lingual capabilities.
Table 4. Summary of recent reasoning-based retrieval benchmarks.
Dataset Year Task Type Reasoning Required Domain
RAR-b (Xiao et al., 2024a) 2024 Reasoning-Intensive Retrieval temporal, numerical, spatial, symbolic General
ImpliRet (Taghavi et al., [n. d.]) 2025 Reasoning-Intensive Retrieval arithmetic, temporal, commonsense General
BRIGHT (Su et al., 2025) 2025 Reasoning-Intensive Retrieval Symbolic, math, code General
InstructIR (Oh et al., 2024) 2024 Instructional IR Instruction following General
InfoSearch (Zhou et al., 2025) 2024 Instructional IR Customized instruction following General
FollowIR (Weller et al., 2024) 2024 Instructional IR Long-Form instruction comprehension General
MAIR (Sun et al., 2024) 2024 Instructional IR Diverse instruction following General
BIRCO (Wang et al., 2024b) 2024 Instructional IR Compositional, Complex Objectives General
mFollowIR (Weller et al., 2025) 2025 Instructional IR Multi-lingual instruction following General
IFIR (Song et al., 2025) 2025 Instructional IR Domain-specific instruction following Biomedical, Legal, Technical
CLERC (Hou et al., 2024) 2024 Case/Legal Retrieval Precedent Reasoning, Legal Logic Legal
Bar Exam/ Housing Statute (Zheng et al., 2025) 2025 Case/Legal Retrieval Symbolic, Legal Reasoning Legal
6.5. Reasoning-aware Embedding
The landscape of IR is rapidly evolving with the emergence of language models capable of complex reasoning and instruction following. Traditional embedding methods, which primarily encode surface-level semantic similarity, often fall short in scenarios requiring logical inference, multi-hop reasoning, or constraint satisfaction. This has motivated the development of reasoning-aware embedding techniques—embedding models explicitly designed to encode latent reasoning paths, logical structures, and task-specific constraints into vector representations.
Two complementary research directions highlight the demand for such embeddings: Reasoning-Intensive Retrieval and Instruction-Following IR. Both emphasize retrieval tasks where shallow matching is insufficient, and embeddings must internalize complex relationships.
Reasoning-Intensive Retrieval involves tasks requiring deep inference over queries and documents. Benchmarks such as RAR-b (Xiao et al., 2024a), BRIGHT (Su et al., 2025), and ImpliRet (Taghavi et al., [n. d.]) evaluate systems on arithmetic reasoning, implicit fact chaining, and temporal logic. In high-stakes domains like law, datasets like CLERC (Hou et al., 2024) and the Legal Reasoning Benchmark (Zheng et al., 2025) demand embeddings that capture legal concepts, precedents, and inferable relationships beyond what surface-level semantics can provide. These challenges necessitate reasoning-aware embedding spaces that can support multi-step deductive or abductive retrieval.
Instruction-Following Retrieval, on the other hand, emphasizes alignment with user intent expressed through natural language instructions. Systems like InstructIR (Oh et al., 2024), MAIR (Sun et al., 2024), and InfoSearch (Zhou et al., 2025) show that instructions often introduce constraints or goals requiring interpretation and reasoning. Benchmarks such as FollowIR, IFIR (Song et al., 2025), and mFollowIR (Weller et al., 2025) evaluate how well models can incorporate instruction semantics into retrieval behavior. Embedding-based approaches here must go beyond vanilla query representation—they must fuse instruction semantics and reasoning signals into a unified vector form, again underscoring the need for reasoning-aware embeddings.
These developments collectively reveal a growing shift from simple semantic encoding toward embedding architectures that encode inferential and procedural knowledge, thereby enhancing retrieval in complex, instruction-rich, or high-reasoning domains.
6.6. Other Domain-Specific Embeddings
Beyond general-purpose language and multi-modal embeddings, there has been growing interest in developing domain-specific embedding models tailored for specialized fields such as finance and chemistry  [^2] Code embeddings focus on programming languages and code semantics across tasks, emphasizing data modality rather than a specific application domain, unlike domain-specific embeddings tailored to fields like finance or chemistry.. These domains pose unique challenges due to their specialized terminology, complex semantics, and domain-specific knowledge requirements. The Finance Massive Text Embedding Benchmark (FinMTEB) (Tang and Yang, 2025) targets the financial domain, covering 64 datasets across seven tasks involving diverse financial text types such as news articles, annual reports, ESG disclosures, and earnings calls in both Chinese and English. FinMTEB evaluations reveal that general-purpose embeddings perform poorly on financial tasks, while domain-adapted models like Fin-E5, trained via persona-based synthetic data, achieve significantly better results. Interestingly, simple Bag-of-Words methods sometimes outperform complex dense embeddings on financial semantic similarity tasks, indicating limitations in current embedding techniques for finance. Similarly, the Chemical Text Embedding Benchmark (ChemTEB) (Kasmaee et al., 2024) addresses the chemical sciences, a domain with specialized linguistic and semantic challenges. ChemTEB evaluates 34 models on chemical literature and data, providing insights into their ability to handle chemical terminology, molecular descriptions, and scientific language. This benchmark highlights the need for domain-aware embedding strategies in chemistry, as generic models often fall short in capturing domain-specific knowledge. Both benchmarks contribute standardized evaluation frameworks that drive the advancement of more accurate and efficient embeddings tailored for expert domains.
@@ 6.6. Other Domain-Specific Embeddings
Table 5. Results on MTEB leaderboard as of July 2025. The models are categorized into two groups based on their embedding dimension (Dim.) and the number of parameters (# Params.).
@@
E5-mistral-7b-instruct (Wang et al., 2024c) 4,096 7B 78.47 50.26 88.34 60.21 56.89 84.63 31.40 66.63
@@
Gemini-embedding (Lee et al., [n. d.]) 3072 - 90.05 59.39 87.7 48.59 64.35 85.29 38.28 73.3
Qwen-embedding-8B (Zhang et al., [n. d.]) 1024 8B 90.43 58.57 87.52 51.56 69.44 88.58 34.83 75.22
Qwen-embedding-4B (Zhang et al., [n. d.]) 2560 4B 89.84 57.51 87.01 50.76 68.46 88.72 34.39 74.60
Qwen-embedding-0.6B (Zhang et al., [n. d.]) 4096 0.6B 85.76 54.05 84.37 48.18 61.83 86.57 33.43 70.70
7. Discussions
@@
Dense vs. Sparse Embedding. While constructing dense embeddings based on LLMs has gained significant attention, some researchers are also investigating the acquisition of sparse embeddings by leveraging lexicon-importance distributions derived from LLMs (Chen et al., 2024; Zhuang et al., 2024; Doshi et al., 2024; Nie et al., 2024). (Zhuang et al., 2024) evaluate various embedding types by directly prompting or fine-tuning LLMs. Their findings reveal that in zero-shot retrieval settings, dense and sparse embeddings perform differently across different benchmarks: for instance, dense embeddings outperform sparse ones on BEIR and TREC-DL 2019, while sparse embeddings show better results on TREC-DL 2020 and MSMARCO. However, in supervised fine-tuning settings, dense retrieval significantly outperforms sparse retrieval on datasets like MSMARCO and TREC-DL 2019/2020 when using models such as Llama3-8B and Llama3-70B. Notably, combining both representation types further enhances performance. (Chen et al., 2024) fine-tune multilingual, multi-granularity text embeddings, integrating both dense and sparse retrieval models based on XLM-RoBERTa. Their results show that dense retrieval notably outperforms sparse retrieval on cross-lingual tasks such as MKQA and MIRACL, and that combining both methods also yields superior results, consistent with the findings reported in (Zhuang et al., 2024).
@@
Last token vs. Mean pooling. Two common methods for obtaining embeddings from a sequence of tokens are: i) mean pooling and ii) the embedding of the last [EOS] token. The former calculates the average of token embeddings, which can potentially obscure key information from phrases. In contrast, [EOS] token embedding is susceptible to recency bias, as it heavily relies on the embedding of the final token in the sequence. (Lee et al., 2024b) conducted experiments comparing several methods for obtaining embeddings from the last layer of LLMs, including the [EOS] token embedding, mean pooling, and their proposed latent attention method. The results demonstrate that mean pooling can always better than [EOS] token embedding based on average MTEB scores, regardless of whether causal attention or bidirectional attention settings are used in the LLMs. Furthermore, their latent attention method enhances embedding capabilities, yielding better results than both the [EOS] token embedding and mean pooling. This research highlights the potential for exploring more advanced designs to improve last-layer embeddings for enhanced performance.
@@
In-context Learning (ICL) for Embedding Model. ICL has emerged as a powerful paradigm for enhancing the performance of LLMs in various generative tasks. It enables LLMs to adapt dynamically to the input context by leveraging real-time examples and instructions during inference. This capability to learn from the surrounding context potentially allows LLMs to produce more accurate and context-sensitive embeddings. Recently, researchers have also explored the effectiveness of ICL for LLM-based representation models. For instance, (Jiang et al., 2023a) first investigated incorporating ICL examples into prompt-based methods and demonstrated their effectiveness for STS tasks. Their approach involved constructing ICL examples by generating word-sentence pairs using ChatGPT to align words with sentence semantics from the STS-B training set, as well as utilizing word-definition pairs from the Oxford dictionary, inspired by DefSent (Tsukagoshi et al., 2021). More recently, (Li et al., 2024d) incorporated few-shot examples on the query side to enhance query embeddings during supervised contrastive training. Despite these few-shot examples being randomly sampled, they still improved the model’s performance. Besides, (Springer et al., 2024) proposed repeating the input twice within the context to obtain sequence embeddings, a technique that can be viewed as a specialized form of ICL, and validated its effectiveness in both training-free and supervised contrastive training settings. These studies collectively indicate that ICL is a promising strategy for enhancing representation learning with LLMs. Nevertheless, exploring more diverse and high-quality contextual information holds potential for further improving the effectiveness of this approach.
Figure 7. The performance trend of models with varying sizes on MTEB benchmark.
@@
Scaling Law of Embedding Models. The question of whether scaling laws hold for embedding models based on LLMS is both significant and intriguing. The effectiveness of embedding models is highly sensitive to several training factors, including model size, data size, and data quality (e.g., diversity and quality of negatives), making it difficult to isolate the effect of each factor independently. (Fang et al., 2024) evaluate the quality of dense retrieval models using a contrastive entropy metric and observe a power law relationship between model performance, model size, and data size across various annotation methods and datasets. However, their experiments were conducted on relatively small-scale models, such as BERT models ranging from 0.5M to 110M parameters. Additionally, (Muennighoff et al., 2022) report the performance of models with varying sizes, such as SGPT (125M, 1.3B, 2.7B, 5.8B) and GTR (Base, Large, XL, XXL), on the MTEB benchmark shown in Figure 7, showing consistent improvements as model size increases, even up to 5.8B parameters. (Zhuang et al., 2024) further test performance by directly prompting LLMs for embeddings, finding that LLaMA3-70B-Instruct outperforms LLaMA3-8B-Instruct across different embedding types. Similarly, the Qwen3-Embedding (Zhang et al., [n. d.]) models demonstrate a positive correlation between scale and performance, improving steadily from 0.6B to 4B and 8B; however, the marginal gain from 4B to 8B is notably smaller, suggesting diminishing returns at higher scales. Taken together, this evidence suggests that the scaling law may also apply to embedding models built on large LMs, although more systematic experiments are needed to rigorously validate this trend across broader configurations.
8. Open Problems